Topic summary

House edge

House edge

Extracted from the Wikipedia article Casino game.

House advantage

In American roulette, there are two "zeroes" (0, 00) and 36 non-zero numbers (18 red and 18 black). This leads to a higher house edge compared to European roulette. The chances of a player, who bets 1 unit on red, winning are 18/38 and his chances of losing 1 unit are 20/38. The player's expected value is EV = (18/38 × 1) + (20/38 × (−1)) = 18/38 − 20/38 = −2/38 = −5.26%. Therefore, the house edge is 5.26%. After 10 spins, betting 1 unit per spin, the average house profit will be 10 × 1 × 5.26% = 0.53 units. European roulette wheels have only one "zero" and therefore the house advantage (ignoring the en prison rule) is equal to 1/37 = 2.7%.